# Bipartite and tripartite entanglement in pure dephasing relativistic spin-boson model

### v1

@@ Abstract
We study non-perturbatively the entanglement generation between two and three emitters in an exactly solvable relativistic variant of the spin-boson model, equivalent to the time-independent formulation of the Unruh-DeWitt detector model. We show that (i) (highly) entangled states of the two emitters require interactions very deep into the light cone, (ii) the mass of the field can generically improve the entanglement generation, (iii) unlike the bipartite case where it is possible to generate close to maximally entangled states via the spin-boson interactions, the generation of genuine tripartite entanglement is non-perturbatively hard even at sufficiently long times. Result (iii), in particular, suggests that probing the multipartite entanglement of a relativistic quantum field non-perturbatively requires either different probe-based techniques or variants of the UDW model. Along the way we provide the regularity conditions for the $N$-emitter model to have well-defined ground states in the Fock space.
@@ I Introduction
The Unruh-DeWitt (UDW) model [1, 2] is a well-known model of light-matter interaction that has been used to study relativistic phenomena predicted from quantum field theory (QFT) in curved spacetimes, such as the Unruh and Hawking effects [3, 4, 5, 6, 1, 2, 7]. The model has been generalized in different ways and applied to many different contexts to study entanglement and communication (see, e.g., [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]). In [29], it was argued that it is sometimes useful to view the UDW model as a relativistic variant of the spin-boson (SB) model [30, 31, 32]. Physically, the main difference between the standard UDW model and the relativistic spin-boson (RSB) model is that the latter is described by a time-independent Hamiltonian — as a closed system — while the former is time-dependent. Terminology-wise, a two-level system is often called a particle detector [^1] This is motivated by Unruh and Hawking effects, as particle number is observer-dependent [1, 3, 33, 34, 35]. in the UDW model, and an emitter in the (R)SB model; we will use them interchangeably.
@@
Here we will first revisit the problem of entanglement generation between two UDW detectors interacting with the vacuum state of the non-interacting scalar field, and then initiate the study of tripartite entanglement generation in this model. We are particularly interested in nonperturbative results by going into the pure dephasing regime, i.e., when the interaction Hamiltonian commutes with the detectors’ free Hamiltonian [27]. The special case of this is when the detectors are gapless, which has been studied in the UDW framework in the context of entanglement harvesting and relativistic communication [20, 22, 36, 26, 37]. We will, however, do this using the closed-system RSB formulation following [29] for several reasons: it gives us a basis for controlling the ultraviolet/infrared (UV/IR) regularity of the model, it is easier to interpret some of the results we present regarding the causal behaviour of the model, and it follows more closely the standard practice in quantum optics. We note that while the dynamics of the pure dephasing regime and the gapless regime are the same, the kinematics are different: for example, gapless RSB model has degenerate joint ground states [29, 30, 31, 32].
@@
More specifically, we use the non-perturbative RSB model to study how the entanglement generated between two and three initially uncorrelated detectors via local interactions with the field depends its wave properties dictated by the strong Huygens principle, which governs the support of the Green’s functions of the wave equation [38, 39, 40]. There are primarily two reasons that motivate our analysis. First, on the one hand there is a no-go theorem forbidding entanglement extraction from the vacuum in the gapless model if the detectors’ interaction regions are spacelike-separated [22], while for massless field in (3+1)D Minkowski spacetime two detectors can be entangled (as expected) if we allow for causal contact [37]. On the other hand, perturbatively the maximal entanglement occurs at the light cone [41] as it is enhanced by signalling, which in turn depends on the strong Huygens principle. Second, we would like to study the tripartite entanglement scenario, which has been done for a closely related delta-coupled variant [28, 42, 43], as this will inform us of the extent the gapless model is useful for studying multipartite entanglement in the context of QFT in curved spacetimes. These will in turn close the (remaining) gap in the literature for gapless UDW model.
@@
In this work, we show that (i) (highly) entangled states of the two emitters require interactions very deep into the light cone — the entanglement cone marking the boundary for nonzero entanglement is deeper in the light cone interior with increasing mass and spacetime dimension; (ii) the mass of the field can generically improve the entanglement generation, at the expense of longer interaction time; we argue, however, that this improvement is somewhat independent of the strong Huygens principle; (iii) unlike the bipartite case where it is possible to generate close to maximally entangled states via the spin-boson interactions, the generation of genuine tripartite entanglement is non-perturbatively hard even at sufficiently long times without any aid from entanglement distillation protocols. Result (iii), in particular, suggests that probing the multipartite entanglement of a relativistic quantum field non-perturbatively via localized probes requires either different probe-based techniques or variants of the UDW model.
@@
This work is organized as follows. In Section II we briefly sketch the setup and establish notation. In Section III we discuss the entanglement dynamics for two detectors/emitters and state some UV/IR regularity requirements for the RSB model. In Section IV we discuss the entanglement dynamics for three detectors/emitters. To ensure that the focus is on the entanglement dynamics, we relegate the remaining technical details regarding the UV/IR regularity of the RSB model to the Appendix A. We adopt the convention that $c=\hbar=1$ and use mostly plus signature for the metric. The symbol $\mathsf{x}$ denotes a spacetime event.
@@ II Setup
In this section, we briefly cover the basic ingredients involved in our analysis. We will first briefly review scalar field theory in the algebraic quantum field theory (AQFT), the formulation of UDW model as a closed-system RSB model, and the strong Huygens principle. In order to focus on the physics, we will only cover the bare minimum required to follow this work. The readers can find much more detailed mathematical construction in the following references: for AQFT, we mainly rely on a very accessible exposition of the scalar field theory in [44] (see also [45, 46, 47]), while for the UDW model we summarize the exposition in [29], which is essentially inspired by [31, 32].
@@ II.1 Scalar field theory in curved spacetime
where $\square=\nabla_{a}\nabla^{a}$ is the d’Alembert operator, $m$ is the mass parameter of the field, $R$ is the Ricci scalar curvature and $\xi\geq 0$ quantifies the degree of non-minimal coupling to gravity. Global hyperbolicity and the nature of the Klein-Gordon differential operator $P$ guarantee that the initial value problem for (1) is well-posed and the spacetime admits a foliation of the form =×$\mathcal{M}=\mathbb{R}\times\Sigma$, where $\Sigma$ is any Cauchy surface that serves as a constant-time slice. For example, we can foliate Minkowski spacetime, the simplest spacetime with zero curvature, in terms of constant-$t$ spacelike hypersurfaces where $t$ is the usual inertial time coordinate.
As a hyperbolic differential operator, $P$ admits unique advanced and retarded propagators $E_{A}(\mathsf{x},\mathsf{x}^{\prime})=E_{R}(\mathsf{x}^{\prime},\mathsf{x})$: they are Green’s functions for $P$ where $E_{R}(\mathsf{x},\mathsf{y})$ [resp. $E_{A}(\mathsf{x},\mathsf{y})$] vanishes if $\mathsf{x}$ is not contained in the causal future (past) of $\mathsf{y}$. We can then define the so-called causal propagator [^2] In [29] it was clarified how the various conventions of the Green’s functions lead to different conventions of the causal propagator, and they can be somewhat confusing. This is because these variations are not fixed solely by the metric signature. Fortunately, in most practical calculations, it does not really matter, as they only affect the overall minus sign that one can typically track. Here we follow [44], which seems to be one of the cleanest.
@@
which is now a kernel of $P$. The causal propagator is important as it appears in the covariant canonical commutation relation (CCR) of the quantized scalar field:
[$𝟙\displaystyle[\hat{\phi}(\mathsf{x}),\hat{\phi}(\mathsf{y})]=\mathsf{i}E(\mathsf{x},\mathsf{y})\openone\,,$ (3)
where $𝟙\openone$ is the identity operator. $E(\mathsf{x},\mathsf{y})$ is a (bi-)distribution since $\hat{\phi}(\mathsf{x})$ is an operator-valued distribution, i.e., it is only a valid operator acting on some Hilbert space after smearing with some appropriate test function $f$. The mathematical formulation based on AQFT thus defines the smeared field operator
@@
In order to do physics, we need to specify states in addition to observables. A state is a $\mathbb{C}$-linear functional $\omega:\mathcal{A}(\mathcal{M})\to\mathbb{C}$ satisfying positivity $\omega(A^{\dagger}A)\geq 0$ and normalization $1\omega(\openone)=1$ for $A\in\mathcal{A}(\mathcal{M})$, i.e., it is a map that takes in an observable $A$ as an input and outputs its expectation value $\omega(A)$. Through the Gelfand-Naimark-Segal (GNS) reconstruction theorem, we can obtain the Hilbert space representation of the theory, with the GNS vector |$\ket{\Omega_{\omega}}$ identified as the “vacuum state” of the theory, and expectation values are given by $\omega(A)=\braket{\Omega_{\omega}}{\pi_{\omega}(A)}{\Omega_{\omega}}$. Here $\pi_{\omega}:\mathcal{A}(\mathcal{M})\to L(\mathcal{H}_{\omega})$ is a GNS representation such that $A$ is represented as a linear operator $\pi_{\omega}(A)\in L(\mathcal{H}_{\omega})$ acting on a (dense subset of) the GNS Hilbert space $\mathcal{H}_{\omega}$. The main advantage of the AQFT framework is that it treats all Hilbert space representations democratically, as there are infinitely many unitarily inequivalent representations of the CCR algebra [46, 47, 45, 44], a feature that is important for studying physical systems that are directly defined in the thermodynamic limit [48]. Indeed, the spin-boson model has been studied in the algebraic framework for the same reason [32, 31, 30, 49, 50].
For physical applications, not all states are physically relevant. First, in a relativistic context, we need the states to be Hadamard states [51, 52, 53]. Concretely, in $(3+1)$-dimensional spacetimes, this corresponds to the states whose Wightman two-point functions have the correct “short-distance (UV) behaviour” and they take the form
@@
Second, we need the GNS Hilbert space $\mathcal{H}_{\omega}$ to be separable: this is a priori not guaranteed (see, e.g., [54] for an explicit example in quantum spin chains in the thermodynamic limit). If we restrict, however, to a subclass of Hadamard states that are quasifree (Gaussian), i.e., those whose Wightman $n$-point functions are fully specified by their 1- and 2-point functions via Wick’s theorem, then the resulting quasifree representations give rise to Fock space $\mathcal{H}_{\omega}=\mathfrak{F}(\mathcal{H})$ (which is separable), where $\mathcal{H}$ [45, 47, 46, 44] is the so-called one-particle Hilbert space associated with the positive-frequency solutions $\varphi_{1},\varphi_{2}$ of the Klein-Gordon equation with Klein-Gordon inner product
@@
Following this construction outlined in [46, 45, 47, 29, 27, 23], we recover the standard Fock space construction in canonical quantization, where we can write the field operator in terms of the Fourier mode decomposition
@@
and the creation and annihilation operators obey the CCR [$𝟙[\hat{a}_{\bm{k}},\hat{a}_{{\bm{k}}^{\prime}}^{\dagger}]=\delta^{n}({\bm{k}}-{\bm{k}}^{\prime})\openone$. The Fock vacuum state of the theory, associated with the GNS vector, is conventionally denoted as |$\ket{0}$ and satisfies the distributional equality $\hat{a}_{\bm{k}}\ket{0}=0$ for all ${\bm{k}}$. All other states with finite particle number can then be built from applications of (smeared) creation operators on the vacuum state and taking linear combinations.
For many practical computations and in order to avoid domain issues associated with unbounded operators $\hat{\phi}(f)$, it is often convenient to work instead with the so-called Weyl algebra $\mathcal{W}(\mathcal{M})$, which is formally an exponentiated version of the field operator, i.e., it is a $C^{*}$-algebra generated by objects of the form $e^{\mathsf{i}\phi(f)}$ (see [45] for more details). One can then find the corresponding state whose GNS construction gives the same physical theory, with the GNS representation $\Pi_{\omega}:\mathcal{W}(\mathcal{M})\to\mathcal{B}(\mathcal{H}_{\omega})$ related via derivatives. [^4] Strictly speaking, the states used for $\mathcal{A}(\mathcal{M})$ and $\mathcal{W}(\mathcal{M})$ are not the same, but there is a bijection between them [55].We note that the state $\omega$ is completely specified by its Wightman $n$-point function:
@@ II.2 UDW model as a time-independent RSB model
The setup we have in mind involves two UDW detectors or emitters A and B interacting locally with a relativistic scalar field. Let us recall the standard prescription based on the covariant formulation of the original UDW model in terms of the Hamiltonian density in the interaction picture [9].
In the original UDW framework, we first assume that the center of mass (COM) of each detector follows a timelike trajectory $\mathsf{z}_{j}(\tau_{j})$, $j\in\{\text{A},\text{B}\}$, parametrized by their own proper time $\tau_{j}$, and then in the COM frame we establish the so-called Fermi normal coordinates (FNC) adapted to the trajectory $\overline{\mathsf{x}}\equiv(\tau,\overline{{\bm{x}}})$ where the COM trajectory is parametrized by $\mathsf{z}_{j}(\tau_{j})\equiv(\tau_{j},\overline{\bm{0}})$ [56]. The point of setting up an FNC is to be able to make a physically reasonable assumption that the interaction between the detector and the field is specified by the spacetime smearing of the form $f(\mathsf{x}(\overline{\mathsf{x}}))=\chi(\tau)F(\overline{{\bm{x}}})$, i.e., in the rest frame of the detector we can tell apart the spatial profile $F(\overline{\bm{x}})$ (associated with, say, atomic orbitals of an atom) from the switching function $\chi(\tau)$ that determines the duration and strength of the interaction. This implicitly assumes that the spatial profile is time-independent within the rest-frame, i.e., we are working in a regime where the detector is (Born-)rigid. The interaction Hamiltonian density in the interaction picture thus takes the form
@@
where $\mathcal{T}_{t}$ is a time-ordering symbol with respect to some global time function. In most cases, one has to perform a perturbative Dyson series truncation, and even then it is mostly intractable except in special cases, such as (conformally) flat spacetimes and specific trajectories, unless one takes a specific limit where the unitary operator above can be calculated exactly — this happens for the so-called delta-coupling and pure-dephasing models [27, 26, 20, 24, 22, 19, 21, 23, 28, 25].
As argued in [29], the UDW model has the nice property that its interaction can be made strictly local in spacetime, so that two detectors A and B can be said to be spacelike separated or not by checking the supports of their interaction profiles $f^{(j)}$. This support region is often taken to represent the portion of spacetime operationally accessible to the corresponding observer, although the algebra of observables that the observer can effectively access can be strictly larger than the interaction region. [^5] For example, a pointlike detector has interaction profile $f$ supported only along its timelike trajectory, but the accessible algebra of observables is prescribed by its timelike envelope, the smallest causal diamond containing $\operatorname{\text{supp}}f$. Taking this difference into account, we can instead reframe the UDW model in the spirit of standard quantum optics, which goes by the name of spin-boson (SB) models. Traditionally, the SB model is defined in flat spacetime and directly in the momentum representation: for a single two-level emitter, its Hamiltonian reads
@@
where $\Omega$ is the energy gap, $\Delta$ measures the shift due to the detector’s internal “magnetic” response, the third term is the field’s free Hamiltonian, and $F_{\bm{k}}$ is called the coupling function that prescribes which modes the detector couples to and by how much. By construction, this model is time-independent as the coupling functions are not dependent on time. We can think of this as a “closed system” formulation of the UDW model where we model the interaction to be always on. Despite this constant-switching interaction, the spacetime locality of observers are still well-defined in the sense that the support of the interaction is localized in space, while the observables accessible to each observer are localized in space and time. Reformulating the UDW model in terms of the constant-switching SB Hamiltonian in (17) we have
@@
with respect to some foliation associated with constant-$\tau$ surfaces. Note that the coupling constants are dimensionful as $[\lambda_{j}]=[\text{Length}]^{(n-3)/2}$ in $(n+1)$-dimensional spacetime.
@@
There are obvious limitations to this formulation that the original UDW model (15) does not have [29]. As currently stated, it only works for static trajectories with respect to the quantization frame, as otherwise $F_{{\bm{k}}}^{(j)}$ in Eq. (20) would be time-dependent, which would correspond to time-dependent coupling (as is the case for Unruh effect or adiabatic switching). It follows that, for multiple detectors, all COM trajectories are static relative to one another, and must align with the Killing vector field generating time translations of the spacetime. Consequently, the SB framework with a time-independent Hamiltonian by default excludes some of the natural settings where the usual UDW approach thrives, e.g., the analysis of relative motion or accelerating trajectories. We will not discuss how to handle these issues in this work, and we refer to [57] for how this is handled in the case of the Unruh effect.
@@
    It allows for a rigorous treatment of the model, i.e., whether it has good UV and IR properties. For example, in the single-emitter case, the existence of ground states of the model depends on the UV/IR regularity of the coupling function $F_{\bm{k}}$ [32, 30, 31]. In the UDW language, this corresponds to making an appropriate choice of spatial smearing $F({\bm{x}})$, whose properties are tied to the Klein-Gordon equation and the spacetime dimension [29]. [^6] In the UDW framework, it is often the case that Gaussian smearing is used for calculational convenience, but its non-compact support leads to mild and quantifiable degrees of causality violation in these models [58, 59, 60, 61]. Provided the detectors are not pointlike, spatial smearings often grant good UV regularity; what remains is to check that the model is also IR regular. There have been a few nice developments providing rigorous studies of the thermalization of an Unruh-DeWitt detector in the weak-coupling regime [57, 62, 63].

### v3

@@ Abstract
We study nonperturbatively the entanglement generation between two and three emitters in an exactly solvable relativistic variant of the spin-boson model, equivalent to the time-independent formulation of the Unruh-DeWitt detector model. We show that (i) (highly) entangled states of the two emitters require interactions very deep into the light cone, (ii) the mass of the field can generically improve the entanglement generation, (iii) while it is possible to find regimes with genuine Greenberger–Horne–Zeilinger-like tripartite entanglement, it is difficult find regimes where tripartite entanglement can be easily shown to be significant or classified. Result (iii), in particular, suggests that probing the multipartite entanglement of a relativistic quantum field nonperturbatively requires either different probe-based techniques or variants of the Unruh-DeWitt model. Along the way, we provide the regularity conditions for the $N$-emitter model to have well-defined ground states in the Fock space.
@@ I Introduction
The Unruh-DeWitt (UDW) model [Unruh1979evaporation, DeWitt1979] is a well-known model of light-matter interaction that has been used to study relativistic phenomena predicted from quantum field theory (QFT) in curved spacetimes, such as the Unruh and Hawking effects [hawking1975particle, wald1994quantum, Crispino2008review, Aubry2014derivative, Unruh1979evaporation, DeWitt1979, tjoa2022unruh]. The model has been generalized in different ways and applied to many different contexts to study entanglement and communication (see, e.g., [Lopp2021deloc, Tales2020GRQO, perche2022localized, Nadine2021delocharvesting, doukas2013unruh, hu2022qhorad, hotta2020duality, perche2023particle, gale2023relativisticCOM, tjoa2023qudit, perche2024udwfromQFT, tjoa2022capacity, Landulfo2016communication, Simidzija2020capacity, Simidzija2018no-go, kasprzak2024transmission, pologomez2024delta, gallock2025qudit, landulfo2024broadcast, tjoa2023nonperturbative, mendez2023tripartite]). In [Tjoa2024interacting], it was argued that it is sometimes useful to view the UDW model as a relativistic variant of the spin-boson (SB) model [amann1991spinboson, spohn1989spinboson, fannes1988equilibrium]. Physically, the main difference between the standard UDW model and the relativistic spin-boson (RSB) model is that the latter is described by a time-independent Hamiltonian—as a closed system—while the former is time-dependent. Terminology-wise, a two-level system is often called a particle detector [^1] This is motivated by Unruh and Hawking effects, as particle number is observer-dependent [Unruh1979evaporation, hawking1975particle, Takagi1986noise, birrell1984quantum, unruhwald1984detector]. in the UDW model, and an emitter in the (R)SB model; we will use them interchangeably.
@@
Here we will first revisit the problem of entanglement generation between two UDW detectors interacting with the vacuum state of the noninteracting scalar field, and then initiate the study of tripartite entanglement generation in this model. We are particularly interested in nonperturbative results by going into the pure dephasing regime, i.e., when the interaction Hamiltonian commutes with the detectors’ free Hamiltonian [tjoa2023nonperturbative]. The special case of this is when the detectors are gapless, which has been studied in the UDW framework in the context of entanglement harvesting and relativistic communication [Landulfo2016communication, Simidzija2018no-go, Simidzija2017coherent, landulfo2024broadcast, perche2024closedform]. We will, however, do this using the closed-system RSB formulation following [Tjoa2024interacting] for several reasons: it gives us a basis for controlling the ultraviolet/infrared (UV/IR) regularity of the model, it makes it easier to interpret some of the results we present regarding the causal behavior of the model, and it follows more closely the standard practice in quantum optics. We note that while the dynamics of the pure dephasing regime and the gapless regime are the same, the kinematics are different: for example, the gapless RSB model has degenerate joint ground states [Tjoa2024interacting, amann1991spinboson, spohn1989spinboson, fannes1988equilibrium].
@@
More specifically, we use the nonperturbative RSB model to study how the entanglement generated between two and three initially uncorrelated detectors via local interactions with the field depends on its wave properties dictated by the strong Huygens principle, which governs the support of the Green’s functions of the wave equation [McLenaghan1974huygens, Sonego1992huygenscurved, Faraoni2019huygens]. There are primarily two reasons that motivate our analysis. First, on the one hand there is a no-go theorem forbidding entanglement extraction from the vacuum in the gapless model if the detectors’ interaction regions are spacelike-separated [Simidzija2018no-go], while for a massless field in (3+1)D Minkowski spacetime two detectors can be entangled (as expected) if we allow for causal contact [perche2024closedform]. On the other hand, perturbatively the maximal entanglement occurs at the light cone [tjoa2021harvesting] as it is enhanced by signalling, which in turn depends on the strong Huygens principle. Second, we would like to study the tripartite entanglement scenario, which has been done for a closely related delta-coupled variant [mendez2023tripartite, mendez2022tripartite, membrere2023tripartite], as this will inform us of the extent to which the gapless model is useful for studying multipartite entanglement in the context of QFT in curved spacetimes. These will in turn close the (remaining) gap in the literature for the gapless UDW model.
@@
In this work, we show that (i) (highly) entangled states of the two emitters require interactions very deep into the light cone—the entanglement cone marking the boundary for nonzero entanglement is deeper in the light cone interior with increasing mass and spacetime dimension; (ii) the mass of the field can generically improve the entanglement generation, at the expense of longer interaction time; we argue, however, that this improvement is somewhat independent of the strong Huygens principle; (iii) while it is possible to find regimes with genuine GHZ-like tripartite entanglement, the parameter space for which this occurs is very small. Furthermore, it is difficult to find regimes where tripartite entanglement can be easily shown to be significant or can be easily classified into different inequivalent types. Result (iii), in particular, suggests that probing the multipartite entanglement of a relativistic quantum field nonperturbatively via localized probes requires either different probe-based techniques or variants of the UDW model.
@@
This work is organized as follows. In Sec. II, we briefly sketch the setup and establish notation and state some UV/IR regularity requirements for the RSB model. In Sec. III we discuss the entanglement dynamics for two detectors/emitters. In Sec. IV we discuss the entanglement dynamics for three detectors/emitters. We adopt the convention that $c=\hbar=1$ and use mostly plus signature for the metric. The symbol $\mathsf{x}$ denotes a spacetime event.
@@ II Setup
In this section, we briefly cover the basic ingredients involved in our analysis. We will first briefly review scalar field theory in the algebraic quantum field theory (AQFT), the formulation of UDW model as a closed-system RSB model, and the strong Huygens principle. In order to focus on the physics, we will only cover the bare minimum required to follow this work. The readers can find much more detailed mathematical construction in the following references: for AQFT, we mainly rely on a very accessible exposition of the scalar field theory in [hack2015cosmological] (see also [fewster2019algebraic, KayWald1991theorems, Khavkhine2015AQFT]), while for the UDW model we summarize the exposition in [Tjoa2024interacting], which is essentially inspired by [spohn1989spinboson, fannes1988equilibrium].
@@ II.1 Scalar field theory in curved spacetime
where $\square=\nabla_{a}\nabla^{a}$ is the d’Alembert operator, $m$ is the mass parameter of the field, $R$ is the Ricci scalar curvature and $\xi\geq 0$ quantifies the degree of nonminimal coupling to gravity. Global hyperbolicity and the nature of the Klein-Gordon differential operator $P$ guarantee that the initial value problem for (1) is well posed and the spacetime admits a foliation of the form $\mathcal{M}=\mathbb{R}\times\Sigma$, where $\Sigma$ is any Cauchy surface that serves as a constant-time slice. For example, we can foliate Minkowski spacetime, the simplest spacetime with zero curvature, in terms of constant-$t$ spacelike hypersurfaces where $t$ is the usual inertial time coordinate.
As a hyperbolic differential operator, $P$ admits unique advanced and retarded propagators $E_{A}(\mathsf{x},\mathsf{x}^{\prime})=E_{R}(\mathsf{x}^{\prime},\mathsf{x})$: they are the Green’s functions for $P$ where $E_{R}(\mathsf{x},\mathsf{y})$ [resp. $E_{A}(\mathsf{x},\mathsf{y})$] vanishes if $\mathsf{x}$ is not contained in the causal future (past) of $\mathsf{y}$. We can then define the so-called causal propagator [^2] In [Tjoa2024interacting] it was clarified how the various conventions of the Green’s functions lead to different conventions of the causal propagator, and they can be somewhat confusing. This is because these variations are not fixed solely by the metric signature. Fortunately, in most practical calculations, it does not really matter, as they only affect the overall minus sign that one can typically track. Here we follow [hack2015cosmological], which seems to be one of the cleanest.
@@
which is now a kernel of $P$. The causal propagator is important as it appears in the covariant canonical commutation relation (CCR) of the quantized scalar field
[…]
@@
[…]
For physical applications, not all states are physically relevant. First, in a relativistic context, we need the states to be Hadamard states [Radzikowski1996microlocal, fewster2013hadamard, fewster2025hadamard]. Concretely, in $(3+1)$-dimensional spacetimes, this corresponds to the states whose Wightman two-point functions have the correct “short-distance (UV) behavior” and they take the form
@@
Second, we need the GNS Hilbert space $\mathcal{H}_{\omega}$ to be separable: this is a priori not guaranteed (see, e.g., [witten2022doesQFTmake] for an explicit example in quantum spin chains in the thermodynamic limit). If we restrict, however, to a subclass of Hadamard states that are quasifree (Gaussian), i.e., those whose Wightman $n$-point functions are fully specified by their 1- and 2-point functions via Wick’s theorem, then the resulting quasifree representations give rise to Fock space $\mathcal{H}_{\omega}=\mathfrak{F}(\mathcal{H})$ (which is separable), where $\mathcal{H}$ [fewster2019algebraic, Khavkhine2015AQFT, KayWald1991theorems, hack2015cosmological] is the so-called one-particle Hilbert space associated with the positive-frequency solutions $\varphi_{1},\varphi_{2}$ of the Klein-Gordon equation with Klein-Gordon inner product
@@
Following this construction outlined in [KayWald1991theorems, fewster2019algebraic, Khavkhine2015AQFT, Tjoa2024interacting, tjoa2023nonperturbative, kasprzak2024transmission], we recover the standard Fock space construction in canonical quantization, where we can write the field operator in terms of the Fourier mode decomposition
@@
[…]
For many practical computations and in order to avoid domain issues associated with unbounded operators $\hat{\phi}(f)$, it is often convenient to work instead with the so-called Weyl algebra $\mathcal{W}(\mathcal{M})$, which is formally an exponentiated version of the field operator, i.e., it is a $C^{*}$-algebra generated by objects of the form $e^{\mathsf{i}\phi(f)}$ (see [fewster2019algebraic] for more details). One can then find the corresponding state whose GNS construction gives the same physical theory, with the GNS representation $\Pi_{\omega}:\mathcal{W}(\mathcal{M})\to\mathcal{B}(\mathcal{H}_{\omega})$ related via derivatives. [^4] Strictly speaking, the states used for $\mathcal{A}(\mathcal{M})$ and $\mathcal{W}(\mathcal{M})$ are not the same, but there is a bijection between them [ruep2021weakly]. We note that the state $\omega$ is completely specified by its Wightman $n$-point function:
@@ II.2 UDW model as a time-independent RSB model
The setup we have in mind involves two UDW detectors or emitters A and B interacting locally with a relativistic scalar field. Let us recall the standard prescription based on the covariant formulation of the original UDW model in terms of the Hamiltonian density in the interaction picture [Tales2020GRQO].
In the original UDW framework, we first assume that the center of mass (COM) of each detector follows a timelike trajectory $\mathsf{z}_{j}(\tau_{j})$, $j\in\{\text{A},\text{B}\}$, parametrized by their own proper time $\tau_{j}$, and then in the COM frame we establish the so-called Fermi normal coordinates (FNC) adapted to the trajectory $\overline{\mathsf{x}}\equiv(\tau,\overline{{\bm{x}}})$ where the COM trajectory is parametrized by $\mathsf{z}_{j}(\tau_{j})\equiv(\tau_{j},\overline{\bm{0}})$ [poisson2011motion]. The point of setting up an FNC is to be able to make a physically reasonable assumption that the interaction between the detector and the field is specified by the spacetime smearing of the form $f(\mathsf{x}(\overline{\mathsf{x}}))=\chi(\tau)F(\overline{{\bm{x}}})$, i.e., in the rest frame of the detector we can tell apart the spatial profile $F(\overline{{\bm{x}}})$ (associated with, say, atomic orbitals of an atom) from the switching function $\chi(\tau)$ that determines the duration and strength of the interaction. This implicitly assumes that the spatial profile is time-independent within the rest frame, i.e., we are working in a regime where the detector is (Born-)rigid. The interaction Hamiltonian density in the interaction picture thus takes the form
@@
where $\mathcal{T}_{t}$ is a time-ordering symbol with respect to some global time function. In most cases, one has to perform a perturbative Dyson series truncation, and even then it is mostly intractable except in special cases, such as (conformally) flat spacetimes and specific trajectories, unless one takes a specific limit where the unitary operator above can be calculated exactly—this happens for the so-called delta-coupling and pure-dephasing models [tjoa2023nonperturbative, landulfo2024broadcast, Landulfo2016communication, pologomez2024delta, Simidzija2018no-go, tjoa2022capacity, Simidzija2020capacity, kasprzak2024transmission, mendez2023tripartite, gallock2025qudit].
As argued in [Tjoa2024interacting], the UDW model has the nice property that its interaction can be made strictly local in spacetime, so that two detectors A and B can be said to be spacelike separated or not by checking the supports of their interaction profiles $f^{(j)}$. This support region is often taken to represent the portion of spacetime operationally accessible to the corresponding observer, although the algebra of observables that the observer can effectively access can be strictly larger than the interaction region. [^5] For example, a pointlike detector has interaction profile $f$ supported only along its timelike trajectory, but the accessible algebra of observables is prescribed by its timelike envelope, the smallest causal diamond containing $\operatorname{\text{supp}}f$. Taking this difference into account, we can instead reframe the UDW model in the spirit of standard quantum optics, which goes by the name of spin-boson (SB) models. Traditionally, the SB model is defined in flat spacetime and directly in the momentum representation: for a single two-level emitter, its Hamiltonian reads
@@
where $\Omega$ is the energy gap, $\Delta$ measures the shift due to the detector’s internal “magnetic” response, the third term is the field’s free Hamiltonian, and $F_{\bm{k}}$ is called the coupling function that prescribes which modes the detector couples to and by how much. By construction, this model is time-independent as the coupling functions are not dependent on time. We can think of this as a “closed system” formulation of the UDW model where we model the interaction to be always on. Despite this constant-switching interaction, the spacetime locality of observers is still well defined in the sense that the support of the interaction is localized in space, while the observables accessible to each observer are localized in space and time. Reformulating the UDW model in terms of the constant-switching SB Hamiltonian in (17), we have
@@
with respect to some foliation associated with constant-$\tau$ surfaces. Note that the coupling constants are dimensionful as $\lambda_{j}]=[\text{Length}]^{(n-3)/2}$ in $(n+1)$-dimensional spacetime.
@@
There are obvious limitations to this formulation that the original UDW model (15) does not have [Tjoa2024interacting]. As currently stated, it only works for static trajectories with respect to the quantization frame, as otherwise $F_{{\bm{k}}}^{(j)}$ in Eq. (20) would be time-dependent, which would correspond to time-dependent coupling (as is the case for Unruh effect or adiabatic switching). It follows that, for multiple detectors, all COM trajectories are static relative to one another, and must align with the Killing vector field generating time translations of the spacetime. Consequently, the SB framework with a time-independent Hamiltonian by default excludes some of the natural settings where the usual UDW approach thrives, e.g., the analysis of relative motion or accelerating trajectories. We will not discuss how to handle these issues in this work, and we refer to [deBievre2006unruh] for how this is handled in the case of the Unruh effect.
@@
    It allows for a rigorous treatment of the model, i.e., whether it has good UV and IR properties. For example, in the single-emitter case, the existence of ground states of the model depends on the UV/IR regularity of the coupling function $F_{\bm{k}}$ [fannes1988equilibrium, amann1991spinboson, spohn1989spinboson]. In the UDW language, this corresponds to making an appropriate choice of spatial smearing $F({\bm{x}})$, whose properties are tied to the Klein-Gordon equation and the spacetime dimension [Tjoa2024interacting]. [^6] In the UDW framework, it is often the case that Gaussian smearing is used for calculational convenience, but its noncompact support leads to mild and quantifiable degrees of causality violation in these models [Causality2015Eduardo, Bruno2020time-ordering, pipo2021causality, maria2023causality]. Provided the detectors are not pointlike, spatial smearings often grant good UV regularity; what remains is to check that the model is also IR regular. There have been a few nice developments providing rigorous studies of the thermalization of an Unruh-DeWitt detector in the weak-coupling regime [deBievre2006unruh, passegger2024disjoint, fewster2016waiting].